OpenStudy (anonymous):

In the expression (x-1) why is f(1/3) undefined?

OpenStudy (anonymous):

Is that the entire question?

OpenStudy (anonymous):

No, (x-1) is what i got after i simplified the expression. and then it said to explain why f(1) = 0, f(0) = -1, and f(-1) = -2, yet f(1/3) is undefined.

OpenStudy (anonymous):

what was your expression?

OpenStudy (anonymous):

anyhoo here are 2 cases where you might have undefined 1/0 sqrt(x) where x<0 ooh and a third one but i doubt it is the case here 0^0

OpenStudy (anonymous):

okkk, i think i get it

OpenStudy (anonymous):

so it could be f(1/3)=1/((1/3)-(1/3))

OpenStudy (jdoe0001):

hmm, what's the original expression?

OpenStudy (anonymous):

3x^2-4x+1 / 3x-1

OpenStudy (anonymous):

i think you got to where the denominator dissapears but it is still a hole there where 3x-1=0 sovling for x you'd get 1/3 This will come up again when you work with limits. Which are the foundation of calculus.

OpenStudy (anonymous):

 here is graph of x-1

OpenStudy (anonymous):

(3x^2-4x+1 )/( 3x-1) Is the same however However  at 1/3 there will be a hole :)